You agree to make 24 deposits of $500 at the beginning of each month into a bank account. At the end of the 24th month, you will have $13,000 in your account. If the bank compounds interest monthly, what nominal annual interest rate will you be earning, that is what is the APR? (If you are using excel the type is: 1)
Answer #1

HI,

We will use excel RATE formula to find out monthly rate first.

here total deposits nper = 24

per month deposit value pmt = -$500

future value fv = $13,000

rate RATE=?

Excel formula for rate = RATE(nper,pmt,pv,fv,type)

since pv is at present 0 so we will put is zero only.

so rate = RATE(24,-500,0,13000,1) = 0.6347%

now this rate is monthly rate.

for annual rate(APR) we will multiply this rate by 12

so APR = 12*0.6347 = 7.62%

Thanks